A worker's total earnings for one day is $100. He received a $20 fixed payment and consumes 14 hours of leisure. What is the hourly wage rate? Show solutions



1
Expert's answer
2022-02-13T12:11:33-0500

TE= FE+VE

Let hourly wage rate be X

100= 20+ X(24-14)

100= 20+ 10X

10X= 80

X= $8